David Notebook: Difference between revisions

Jump to navigation Jump to search
 
==Daily Log==
'''11/19/14'''
Once again, 1555 locked at 0.05 atm with minimal adjustment. This is a very good sign. I've got a new procedure for getting 633 back if the 780 gets misaligned. No cameras or IR viewers required and it's quicker and easier--put a fair amount of gas in (maybe 1 V or so) so that there aren't many spacial modes without 633 and there's a lot of it if alignment is good. On the 1064 side of the cavity, turn down the power and use a card to overlap the 1064 and 780. Then turn the power back up, lock the 1064, and on the 780 side of the cavity use the other alignment mirror to overlap the transmitted 1064 with the 780. Repeating this 2-3 times will get 633 back depending on how badly aligned it was. Optimal alignment will change slightly when you switch back to whatever pressure you were working at, but it should still easily be detectable and so is not hard to fix.
 
I'm getting closer to being able to do the 2 photon experiment. Both lasers have been locking pretty well at this pressure, 633 is easily visible with just the 1064. Now I just need to practice tuning the laser frequencies and figure out how to keep them that way even when they lock and jump frequencies.
 
I think pre-stabilizing them with reference cavities would help a lot, but I'll keep trying this for a bit and see if I can make it happen without them. There's still probably a few weeks before the Moglabs wavemeter is ready, and the pulsed laser is still not set up either, so it's not really holding much up.
 
'''11/18/14'''
Hot damn! 0.05 atm might be the magic pressure. 1555 locked again today with only small adjustments. Good settings seem to be integrator at 5 kHz, differentiator at 500 kHz (this definitely seems the best setting for all pressures). Differentiator gain seems to give a trade off between power and stability. Prop. gain around 3.75. Optical feedback at 1.4 V (I didn't realize the scale on the oscilloscope had changed--I've normally been doing around 400 mV, but this actually worked well with different gain settings. Maybe more optical feedback, less gain is better if I can get it to work).

Navigation menu